OpenStudy (anonymous):

If the average test score of four students is 85, which of the following could the fifth student score in order to have the new average be greater than 84 and less than 86?

OpenStudy (anonymous):

total = 85*4 = 340 340 +x/5 = 85 340+x +425 x= 125

OpenStudy (anonymous):

sorry thats wrong

hero (hero):

85

hero (hero):

you set it up wrong

OpenStudy (anonymous):

85 is not the correct answer either

OpenStudy (anonymous):

425-340= 85 not 125 :P my bad

OpenStudy (anonymous):

it has to be

OpenStudy (anonymous):

If the average test score of four students is 85, which of the following could the fifth student score in order to have the new average be greater than 84 and less than 86?

OpenStudy (anonymous):

for some reason that's not one of the options given as an answer

OpenStudy (anonymous):

what are the other options

hero (hero):

I think it would be smart to post the options

OpenStudy (anonymous):

A. 70 B. 90 C. 88 D. 83 E. Both C and D

hero (hero):

The reason why you have to post the answer choices with this one is because 85 isn't incorrect, however, there are a range of other numbers that could work. The answer is most likely 83 because it is the closest to 85

hero (hero):

And yes, because if the fifth person scores 83, the average is 84.5 which works.

OpenStudy (anonymous):

Thanks! :)

hero (hero):

But 88 also works so the actual answer is E